OpenStudy (anonymous):

help? find the inverse of the given function and use composition to justify the results of: g(x)= x-3/x

hartnn (hartnn):

could you find inverse? ? let you find inverse function as \(f(x)\) composition means \(f(g(x))\) if your inverse is correct, then \(f(g(x)) = x\)

OpenStudy (anonymous):

would the inverse be, x(x+3)

hartnn (hartnn):

its (x-3)/x or x - 3/x ?

OpenStudy (anonymous):

(x-3)/x sorrry for not being clear

hartnn (hartnn):

y = (x-3)/x xy = x-3 xy - x = -3 x (y-1) = -3 x= -3/(y-1) so your inverse function is -3/(x-1) did u get how i got the inverse ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about using composition to justify the result?

hartnn (hartnn):

if \(f(g(x))=x\) then f and g are inverses of each other g(x) = (x-3)/x f(x) = -3/(x-1)

hartnn (hartnn):

to get f(g(x)) , plug in the expression of g(x) in f(x)

hartnn (hartnn):

wherever you see 'x' in f(x) just replace it by (x-3)/x

OpenStudy (anonymous):

how would u simplify

OpenStudy (anonymous):